OpenStudy (anonymous):

I need help with a question I had a few hours ago on a test. Using the method of undetermined coefficients find the solution to y'' - y' = 4te^(2t) y(0)=0 y'(0)=0

OpenStudy (nubeer):

ok have you solved the left side?

OpenStudy (anonymous):

that's what the problem is asking. The left side is the differential equation.

OpenStudy (nubeer):

Y=Ae^(mx) this is a general form find Y', Y'' and plug them in your equation.. and remember when u r working on left side, the right side is equal to 0 so Y''-Y' = 0 your equation now should be m^2-m = 0 please solve and see if u get this or not.

OpenStudy (anonymous):

yeah so y''-y'=0 \[r^2-r = 0\]\[r(r-1)=0\]\[r _{1} = 0\]\[r _{2} = 1\]\[y_{c}=C _{1}e^0 + C _{2}e^t\]\[y_{c}=C_{1} + C_{2}e^t\]Carrying on from that: \[y _{c} = Ae ^{2t}(Bt+C)\] I think.. that's where I got confused.

OpenStudy (nubeer):

ya its yp now find yp ' and yp'' ... now plug this derivative in equation so yp''-yp' = Ae^(2t)(Bt+C)

OpenStudy (anonymous):

ok cool thanks. Yeah i meant to type yp but I didn't. thanks I got it now.

OpenStudy (nubeer):

ok good :)